OpenStudy (anonymous):

I really don't understand how to factor and solve this x^3-27=0

OpenStudy (p0sitr0n):

x^3-3^3=0

OpenStudy (triciaal):

27 can be written as 3^3 x^3 = 3^3 so x = 3

OpenStudy (acxbox22):

(x-3)^3=0 x-3=0 x=3

OpenStudy (triciaal):

sorry to factor use the difference of 2 cubes x^3 - 3^3 = (x-3)(x-3x^2 *3 + 3 x*3^2